OpenStudy (lifeisadangerousgame):

uppose that hydrogen atoms are prepared in the n = 2 energy level. Calculate the wavelength of light emitted by these excited atoms when returning to the ground state. ANy idea how to do this?

OpenStudy (joannablackwelder):

I would use this formula: \[1/\lambda = R _{H }\left( \frac{ 1 }{ n _{1}^{2} } - \frac{ 1 }{ n _{2} ^{2}}\right)\]

OpenStudy (lifeisadangerousgame):

I'm not given lamba though?

OpenStudy (lifeisadangerousgame):

Ah wait. I understand.

OpenStudy (joannablackwelder):

RH is 1.09678 x 10^-2 nm^-1

OpenStudy (joannablackwelder):

Yeah, lambda is what you are looking for. :-)

OpenStudy (lifeisadangerousgame):

Okay, so n(sub)1 is 2 and n (sub) 2 is 1 right?

OpenStudy (joannablackwelder):

Yep.

OpenStudy (lifeisadangerousgame):

Okay, thank you! Could you help me on another one?

OpenStudy (lifeisadangerousgame):

I'll make another question so you can get a medal

OpenStudy (joannablackwelder):

You're welcome. :-) And I can try.

OpenStudy (lifeisadangerousgame):

Hey for this one, I got 1/1.6 x 10^-18 J, is there anyway I can verify if it's correct? If not, you don't need to trouble yourself with working the problem out c:

OpenStudy (lifeisadangerousgame):

I got 1/-1.6 x 10^-18J sorry